I have a problem for Hw that i dont understand.
the problem says to write the equation in standard form.
heres the equation: 4x^2=12
how would you write that equation in standard form?
thanks for the help!

Answer by palanisamy(496)

The given equation is 4x^2 = 12
4x^2 = 4*3
x^x = 3
x^2-3 = 0